Tuesday, July 3, 2012

Day 37, 7/3/2012 -- Not awesome

Today was kind of a bust in terms of getting any real data. Last night after dinner, Seth and I went back to the high bay to do a simple fix on one of the "general housekeeping" boards in the system that lives in one of the bolometer power crates. Turns out that after we reconnected everything, the cryostat temperature readouts stopped working. And I was informed of this at approximately 1AM. So I went back to the high bay to try and fix it, was thoroughly unsuccessful, Seth tried to debug the issue in the morning before we got in, and we finally tracked it down to connecting mislableled connectors on the power box the way they were indicated -- i.e. incorrectly!

After fixing the labels, we then set in motion a new 'schedule file' for the bolometer system that would automatically put the detector array into a nearly-operational state. Of course, this didn't work completely initially, leading to some time spent debugging. And then the focal plane was taking forever to cool...partially because one of the fridges ran out. Even so, we pushed ahead and tried to get some data as the detectors were still quite cold...but we ended up not being able to get any usable calibration data today. Major bummer. On the plus side, Jeff was able to make some progress on the data analysis for our calibration scans giving us some more confidence that our data is of good quality.

The gondola team continues to make progress; now that the Bemco testing is done they have started to reassemble the entire gondola. A continuing discussion with CSBF revolves around placement and protection of their SIP LDB electronics package, and I believe they are converging on a solution.

Pictures:
https://picasaweb.google.com/104253244018605213307/EBEXInPalestine20120703

No comments:

Post a Comment